# Bilinear Subspace Variational Bayesian Inference for Joint Scattering Environment Sensing and Data Recovery in ISAC Systems

### v1

@@ IV-C3 Convergence Analysis
Despite the above approximations in the calculation of the posterior mean $\boldsymbol{\mu}_{\alpha}=\mathbf{W}_{\alpha}^{-1}\boldsymbol{b}_{\alpha}$, we can show that the proposed BiSVBI still converges to a stationary point of $\mathcal{\mathscr{A}}_{\mathrm{VBI}}$, as stated in the following theorem, whose proof can be found in Appendix -B.
@@ IV-D Turbo-BiSVBI-M Step
$=$ $-\sum_{t_{p}=1}^{T_{p}}\left\langle\kappa\right\rangle\left\|\boldsymbol{y}_{t_{p}}^{p}-\mathbf{U}_{t_{p}}\boldsymbol{\Phi}\left(\boldsymbol{\xi}\right)\boldsymbol{\mu}_{\alpha}^{\left(t\right)}\right\|^{2}$
$-\sum_{t_{d}=1}^{T_{d}}\left\langle\kappa\right\rangle\left\|\boldsymbol{y}_{t_{d}}^{d}-\left\langle\mathbf{X}_{t_{d}}\right\rangle\boldsymbol{\Phi}\left(\boldsymbol{\xi}\right)\boldsymbol{\mu}_{\alpha}^{\left(t\right)}\right\|^{2}$
@@
When $q^{\left(t\right)}\left(\overline{\boldsymbol{v}}\right)=p\left(\tilde{\boldsymbol{v}}\mid\boldsymbol{y};\boldsymbol{\xi}^{\left(t\right)}\right)$, it can be shown that the gradient of $Q\left(\boldsymbol{\xi};\boldsymbol{\xi}^{\left(t\right)}\right)$ is consistent with $\ln p\left(\boldsymbol{y},\boldsymbol{\xi}\right)$ [20], and thus we can apply the gradient ascent method to update $\boldsymbol{\xi}$ as
@@
9:        Update $q^{\left(t\right)}\left(\kappa\right)$ using (38).
@@
11:     Calculate the MMSE estimates $\boldsymbol{\alpha}^{\left(t\right)}$, $\boldsymbol{x}^{\left(t\right)}$, and $\kappa^{\left(t\right)}$, and then update the support set $\hat{\mathcal{S}}$ using (32).
@@ V-B SCVBI to Refine the Angles
where $\mathcal{N}_{k}^{p}=\cup_{t_{p}}\mathcal{N}_{t_{p},k}^{p}$, $p\left(\tilde{\boldsymbol{s}}_{k}\right)=\prod_{l=1}^{\tilde{L}_{k}}\left(\tilde{\lambda}_{k,l}\right)^{\tilde{s}_{k,l}}\left(1-\tilde{\lambda}_{k,l}\right)^{1-\tilde{s}_{k,l}},$and $\tilde{\lambda}_{k,l}$ is the probability of $\tilde{s}_{k,l}=1$, which is set to be $\tilde{\lambda}_{k,l}=\frac{1}{\left|\mathcal{U}_{k}\right|}$ in the simulations. $p\left(\tilde{\boldsymbol{\rho}}_{k,n}\mid\tilde{\boldsymbol{s}}_{k}\right)$ is a conditional Gamma distribution such that the mean value of $\tilde{\rho}_{k,n,l}$ is set to be $O\left(1\right)$ when $\tilde{s}_{k,l}=1$ and a large value when $\tilde{s}_{k,l}=0$. Finally, $p\left(\tilde{\alpha}_{k,n,l}\mid\tilde{\rho}_{k,n,l}\right)=\mathcal{CN}\left(\tilde{\alpha}_{k,n,l};0,\tilde{\rho}_{k,n,l}^{-1}\right),\forall l$.
@@ VII Conclusion
This paper propose a novel joint sensing and data recovery algorithm for massive MIMO-OFDM ISAC systems. A 3D location-domain sparse prior model is firstly introduced to capture the joint sparsity of the MU channel. Then, the joint problem is formulated as a bilinear structured sparse recovery problem with a dynamic position grid and imperfect parameters. An EM-Turbo-BiSVBI algorithm is proposed to solve this joint sensing and data recovery problem, where the E-step performs Bayesian estimation of the location-domain sparse MU channel by exploiting the joint sparsity, and the M-step refines the dynamic position grid and learns the imperfect factors via gradient update. In particular, a BiSVBI algorithm is proposed to solve the bilinear sparse inference problem in the E-step with both theoretical convergence guarantee and low complexity. In addition, an SMUSIC-SCVBI algorithm is proposed in the coarse estimation stage to narrow down the search range and reduce the effective dimension of the location-domain channel, which further reduces the complexity caused by the ultra-high-dimensional sparse channel in the entire 3D space. The proposed algorithm is shown by simulations to achieve much better trade-off between performance and complexity than various baseline algorithms.
@@
Given $q\left(\boldsymbol{x}\right),q\left(\boldsymbol{\rho}\right),q\left(\boldsymbol{s}\right),q\left(\kappa\right)$, the KLD can be calculated as
@@
$-\left\langle\ln p\left(\boldsymbol{y}\mid\boldsymbol{\alpha},\boldsymbol{x},\kappa\right)\right\rangle_{q\left(\boldsymbol{\alpha}\right)q\left(\boldsymbol{x}\right)q\left(\kappa\right)}+C_{1},$
@@
$+\left\langle\kappa\right\rangle\left\langle\left\|\boldsymbol{y}-\mathbf{A}_{\alpha}\boldsymbol{\alpha}\right\|^{2}\right\rangle_{q\left(\boldsymbol{\alpha}\right)q\left(\boldsymbol{x}\right)}+C_{1},$ (54)
@@
Similarly, given $q\left(\boldsymbol{\alpha}\right),q\left(\boldsymbol{\rho}\right),q\left(\boldsymbol{s}\right),q\left(\kappa\right)$, the KLD can be calculated as
@@
$-\left\langle\ln p\left(\boldsymbol{y}\mid\boldsymbol{\alpha},\boldsymbol{x},\kappa\right)\right\rangle_{q\left(\boldsymbol{\alpha}\right)q\left(\boldsymbol{x}\right)q\left(\kappa\right)}+C_{2},$
@@
$\left.+\left\langle\kappa\right\rangle\left\langle\left\|\boldsymbol{y}_{t_{d},n}^{d}-\mathbf{\Psi}_{n}\boldsymbol{x}_{t_{d},n}\right\|^{2}\right\rangle_{q\left(\boldsymbol{\alpha}\right)q\left(\boldsymbol{x}_{t_{d},n}\right)}\right]+C_{2},$ (56)

### v2

@@ IV-C3 Convergence Analysis
Despite the above approximations in the calculation of the posterior mean $\boldsymbol{\mu}_{\alpha}=\mathbf{W}_{\alpha}^{-1}\boldsymbol{b}_{\alpha}$, we can show that the proposed BiSVBI still converges to a stationary point of $\mathcal{\mathscr{A}}_{\mathrm{VBI}}$, as stated in the following theorem, whose proof can be found in Appendix B.
@@ IV-D Turbo-BiSVBI-M Step
$=$ $-\sum_{t_{p}=1}^{T_{p}}\left\langle\gamma\right\rangle\left\|\boldsymbol{y}_{t_{p}}^{p}-\mathbf{U}_{t_{p}}\boldsymbol{\Phi}\left(\boldsymbol{\xi}\right)\boldsymbol{\mu}_{\alpha}^{\left(t\right)}\right\|^{2}$
$-\sum_{t_{d}=1}^{T_{d}}\left\langle\gamma\right\rangle\left\|\boldsymbol{y}_{t_{d}}^{d}-\left\langle\mathbf{X}_{t_{d}}\right\rangle\boldsymbol{\Phi}\left(\boldsymbol{\xi}\right)\boldsymbol{\mu}_{\alpha}^{\left(t\right)}\right\|^{2}$
@@
When $q^{\left(t\right)}\left(\overline{\boldsymbol{v}}\right)=p\left(\overline{\boldsymbol{v}}\mid\boldsymbol{y};\boldsymbol{\xi}^{\left(t\right)}\right)$, it can be shown that the gradient of $Q\left(\boldsymbol{\xi};\boldsymbol{\xi}^{\left(t\right)}\right)$ is consistent with $\ln p\left(\boldsymbol{y},\boldsymbol{\xi}\right)$ [19], and thus we can apply the gradient ascent method to update $\boldsymbol{\xi}$ as
@@
9:        Update $q^{\left(t\right)}\left(\gamma\right)$ using (38).
@@
11:     Calculate the MMSE estimates $\boldsymbol{\alpha}^{\left(t\right)}$, $\boldsymbol{x}^{\left(t\right)}$, and $\gamma^{\left(t\right)}$, and then update the support set $\hat{\mathcal{S}}$ using (32).
@@ V-B SCVBI to Refine the Angles
where $\mathcal{N}_{k}^{p}=\cup_{t_{p}}\mathcal{N}_{t_{p},k}^{p}$, $p\left(\tilde{\boldsymbol{s}}_{k}\right)=\prod_{l=1}^{\tilde{L}_{k}}\left(\tilde{\lambda}_{k,l}\right)^{\tilde{s}_{k,l}}\left(1-\tilde{\lambda}_{k,l}\right)^{1-\tilde{s}_{k,l}},$and $\tilde{\lambda}_{k,l}$ is the probability of $\tilde{s}_{k,l}=1$, which is set to be $\tilde{\lambda}_{k,l}=\frac{1}{\left|\mathcal{U}_{k}\right|}$ in the simulations. $p\left(\tilde{\boldsymbol{\rho}}_{k,n}\mid\tilde{\boldsymbol{s}}_{k}\right)$ is a conditional Gamma distribution such that the mean value of $\tilde{\rho}_{k,n,l}$ is set to be $\mathcal{O}\left(1\right)$ when $\tilde{s}_{k,l}=1$ and a large value when $\tilde{s}_{k,l}=0$. Finally, $p\left(\tilde{\alpha}_{k,n,l}\mid\tilde{\rho}_{k,n,l}\right)=\mathcal{CN}\left(\tilde{\alpha}_{k,n,l};0,\tilde{\rho}_{k,n,l}^{-1}\right),\forall l$.
@@ VII Conclusion
This paper propose a novel joint sensing and data recovery algorithm for massive MIMO-OFDM ISAC systems. A 3D location-domain sparse prior model is firstly introduced to capture the joint sparsity of the MU channel. Then, the joint problem is formulated as a bilinear structured sparse recovery problem with a dynamic position grid and imperfect parameters. An EM-Turbo-BiSVBI algorithm is proposed to solve this joint sensing and data recovery problem, where the E-step performs Bayesian estimation of the location-domain sparse MU channel by exploiting the joint sparsity, and the M-step refines the dynamic position grid and learns the imperfect factors via gradient update. In particular, a BiSVBI algorithm is proposed to solve the bilinear sparse inference problem in the E-step with both theoretical convergence guarantee and low complexity. In addition, a SMUSIC-SCVBI algorithm is proposed in the coarse estimation stage to narrow down the search range and reduce the effective dimension of the location-domain channel, which further reduces the complexity caused by the ultra-high-dimensional sparse channel in the entire 3D space. The proposed algorithm is shown by simulations to achieve much better trade-off between performance and complexity than various baseline algorithms.
@@
Given $q\left(\boldsymbol{x}\right),q\left(\boldsymbol{\rho}\right),q\left(\boldsymbol{s}\right),q\left(\gamma\right)$, the KLD can be calculated as
@@
$-\left\langle\ln p\left(\boldsymbol{y}\mid\boldsymbol{\alpha},\boldsymbol{x},\gamma\right)\right\rangle_{q\left(\boldsymbol{\alpha}\right)q\left(\boldsymbol{x}\right)q\left(\gamma\right)}+C_{1},$
@@
$+\left\langle\gamma\right\rangle\left\langle\left\|\boldsymbol{y}-\mathbf{A}_{\alpha}\boldsymbol{\alpha}\right\|^{2}\right\rangle_{q\left(\boldsymbol{\alpha}\right)q\left(\boldsymbol{x}\right)}+C_{1},$ (54)
@@
Similarly, given $q\left(\boldsymbol{\alpha}\right),q\left(\boldsymbol{\rho}\right),q\left(\boldsymbol{s}\right),q\left(\gamma\right)$, the KLD can be calculated as
@@
$-\left\langle\ln p\left(\boldsymbol{y}\mid\boldsymbol{\alpha},\boldsymbol{x},\gamma\right)\right\rangle_{q\left(\boldsymbol{\alpha}\right)q\left(\boldsymbol{x}\right)q\left(\gamma\right)}+C_{2},$
@@
$\left.+\left\langle\gamma\right\rangle\left\langle\left\|\boldsymbol{y}_{t_{d},n}^{d}-\mathbf{\Psi}_{n}\boldsymbol{x}_{t_{d},n}\right\|^{2}\right\rangle_{q\left(\boldsymbol{\alpha}\right)q\left(\boldsymbol{x}_{t_{d},n}\right)}\right]+C_{2},$ (56)